# The Value of Information in Human-AI Decision-making

### v5

@@ Results.
Figure 1 confirms the expected ranking of AI models by $\operatorname{ACIV}$. AI1, which had access to the two human-uninterpretable features, achieves higher $\operatorname{ACIV}$ than AI2, which only had access to the four human-interpretable features. Note the similar predictive performance of the two AI models.
The posterior predictions of human-AI team performance from our regression model also align with expectations based on the complementary information each model provides (Figure 1 far right). Collapsing across explanation conditions, AI1 (higher ACIV) reduces APE by more than AI2 ($6.27\%$$5.91\%,$$6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively). Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively) [^5] See Appendix D for the significant test results.
Figure 1 also illustrates the effect of the different explanation conditions on human-AI team performance. When the AI has sufficient complementing information (AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation).
@@ 6.1 Chest X-ray Diagnosis
Figure 2(a) shows all the models offer complementary information value to the radiologists’ reports (green distributions improve over the purple distribution), and in the other direction, the radiologists’ reports offer complementary information value to the models (green distributions improve over the orange distributions).
@@
Figure 2(a) shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We further assess the robustness of VisionTransformer’s superiority over the other AI models across many possible payoff functions to test if there is a Blackwell ordering of models in Appendix F and Figure 6.
@@
Figure 2: Demonstration results for both tasks. Distributions show information values. For a), we plot ACIV of radiologist decisions, different choices of AI models and signals that combine AI predictions and radiologist decisions. For b), we plot the information value of the combination of video-level features and agent decisions (including human decisions, AI predictions, and human-AI teams’ decisions). These two figures demonstrate that our framework can be used to a) reveal the complementary information between AI predictions and human decisions and b) analyze how humans use information in data features differently from AI.
As shown in this demonstration, doctors may use our framework to learn how much complementary information value the AI models offer over their decisions, and which model offers the most. [^6] For space constraints, we refer to the full descriptions and results in Appendices F and E.
@@ 6.2 Deepfake Detection
We study a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022). We select the model in the Deepfake Detection Challenge (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data. We use the human decisions and the human-AI team’s decisions from Groh et al. (2022), who collected judgments on the videos from n=5,524 participants recruited on Prolific. We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from Groh et al. (2022)’s use of mean absolute error, but again we prefer the quadratic score because it is a proper scoring rule where truthfully reporting beliefs maximizes the score. Groh et al. (2022) manually label seven video-level features, which we use as binary indicators in place of Algorithm 1 in light of the high dimensional signals: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin. We show the ACIVs of the seven features over the decisions, the AI predictions, and the human-AI decisions in Figure 2(b).
@@
We first compare the information value of the AI predictions to that of the human decisions. Figure 2(b) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$).
@@
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different video-level features over different agents. As shown on the fifth row in Figure 2(b), the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 2(b), the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions.
@@ 7 Discussion and Limitations
Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al. (2023, 2024)). While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments. Though most of our definitions and analysis are focused on decision tasks with single well-defined payoff function, the framework can be readily extended to more complex decision tasks with an ambiguous set of payoff functions or totally unidentifiable payoff functions. We present a robustness analysis framework in Appendix B to calculate the Blackwell ordering of models over proper scoring rules. A demonstration of the robustness analysis is shown in Appendix F.
@@ Appendix A The Combinatorial Nature of the Value of Signals
When decision-makers are provided with multiple signals, the signals have the combinatorial property by nature. Acknowledgd by recent works in decision theory and game theory [Chen and Waggoner, 2016], one signal may have no information value by itself, but it can be complementary to other signals to provide information value. For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ are uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$. In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff. Though we did not explicitly observe the complementation between signals in our survey of human-AI decision-making tasks (see the results of deepfake detection in supplementary materials), we want to note that our framework can be extended to consider this complementation between signals. We use the Shapley value [Shapley, 1953] to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal. $\phi$ is the average of the marginal contribution of a signal in every combination with other signals.
@@ Appendix B Robust Analysis of Information Order
Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem. However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule. In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients [Mclaughlin and Spiess, 2023]. Blackwell’s comparison of signals [Blackwell et al., 1951] is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [Li et al., 2022, Kleinberg et al., 2023].
@@
Proposition B.3 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$. Proposition B.3 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
@@
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix E.
@@ Appendix C Observational Study on the Effect of ACIV on Human-AI Team Performance
In this section, we present the results of an observational study on the effect of $\operatorname{ACIV}$ on human-AI team performance using the dataset from Vodrahalli et al. [2022].
@@
We experiment with the publicly available Human-AI Interactions dataset [Vodrahalli et al., 2022]. The dataset comprises 34,783 unique predictions from 1,088 different human participants on four different binary prediction tasks (“Art”, “Sarcasm”, “Cities” and “Census”). In each of the tasks, human participants provide confidence values about their predictions before ($d^{H}$) and after ($d^{H+AI}$) receiving AI advice from a classifier in the form of the classifier’s confidence values ($d^{AI}$). $d^{H}$, $d^{H+AI}$ and $d^{AI}$ are in the range of $[0,1]$.
@@ Results.
We present the $\operatorname{ACIV}$ of the AI confidence values ($d^{AI}$) over human decisions ($d^{H}$) measured by the payoff function as the mean squared error (MSE) in Figure 3, and show the reduction of the mean squared error (MSE) of the human-AI team over the human-alone baseline in Table 1. We find that between these four tasks, the $\operatorname{ACIV}$ of the AI over human decisions predicts the improvement of the human-AI team performance over the human-alone baseline.
@@ Appendix D Statistical Tests on Experiment Results
where $s_{1}$ and $s_{2}$ are the standard deviations of the two samples, and $N_{1}$ and $N_{2}$ are the sizes of the two samples (see sample sizes and standard deviations in Table 3). We used Bonferroni correction as the multiple comparisons correction to control the overall type I error rate. The results are shown in Table 2.
Table 2: 95% confidence intervals and significance test results (p-values) for the reduction of APE with different AI models and explanations in Figure 1.
P(ILIV-SHAP + SHAP
¿ SHAP)
P(ILIV-SHAP + SHAP
¿ No explanation)
P(AI1 ¿ AI2)
P(AI1 + ILIV-SHAP + SHAP
¿ AI2 + ILIV-SHAP + SHAP)
Reduction of APE with 95% confidence intervals
@@
$6.24\%$
$5.79\%,6.68\%]$
$6.29\%$
$5.80\%,6.78\%]$
$5.47\%$
$4.95\%,5.98\%]$
0.138 / 0.159
@@
Appendix E Demonstration I: Model Comparison on Chest Radiograph Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task [Rajpurkar et al., 2018, Tang et al., 2020, Shreekumar, 2025]. We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
@@ Appendix E Demonstration I: Model Comparison on Chest Radiograph Diagnosis
E.1 Data and Model
We use data from the MIMIC dataset [Goldberger et al., 2000], which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database [Johnson et al., 2019] merged with patient and visit information from the broader MIMIC-IV database [Johnson et al., 2023]. The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. [2019] and Heidenreich et al. [2022]. We label the radiology reports by a rule-based tool [Irvin et al., 2019] and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports. The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$. We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer [Alexey, 2020], SwinTransformer [Liu et al., 2021], ResNet [He et al., 2016], Inception-v3 [Szegedy et al., 2016], and DenseNet [Huang et al., 2017]. Our training set contains 12,228 images, and the validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models is: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, and VisionTransformer with $0.80$.

### v6

@@ Results.
Table˜1 confirms the expected ranking of AI models by $\operatorname{ACIV}$. AI1, which had access to the two human-uninterpretable features, achieves higher $\operatorname{ACIV}$ than AI2, which only had access to the four human-interpretable features. Note the similar predictive performance of the two AIs.
The posterior predictions of human-AI team performance from our regression model also align with expectations based on the complementary information each model provides (Figure˜2 far right). Collapsing across explanation conditions, AI1 (higher ACIV) reduces APE by more than AI2 ($6.24\%$$5.99\%,$$6.5\%]$ vs $5.96\%[5.68\%,6.24\%]$ respectively). Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($6.94\%[6.50\%,7.38\%]$ vs $5.31\%[4.80\%,5.83\%]$ respectively) [^8] See Appendix D for the significant test results.
Figure˜2 also illustrates the effect of the different explanation conditions on human-AI team performance. When the AI has sufficient complementing information (AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($6.94\%[6.50\%,7.38\%]$ for ILIV-SHAP, $5.88\%[5.47\%,6.28\%]$ for SHAP, and $5.96\%[5.50\%,6.42\%]$ for no explanation).
@@ 6.1 Chest X-ray Diagnosis
Figure˜3(a) shows all the models offer complementary information value to the radiologists’ reports (green distributions improve over the purple distribution), and in the other direction, the radiologists’ reports offer complementary information value to the models (green distributions improve over the orange distributions).
@@
Figure˜3(a) shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We further assess the robustness of VisionTransformer’s superiority over the other AI models across many possible payoff functions to test if there is a Blackwell ordering of models in Appendix˜G and Figure˜7.
@@
Figure 3: Distributions show information values. For a), we plot ACIV of radiologist decisions, different choices of AI models and signals that combine AI predictions and radiologist decisions. For b), we plot the information value of the combination of video-level features and agent decisions (including human decisions, AI predictions, and human-AI teams’ decisions). Full information represents all the information available to the human decision-makers–the radiology images for the chest X-ray diagnosis task and the seven video-level features, human decisions, and AI predictions for the deepfake detection task. See Appendix˜H for the sensitivity analysis of Algorithm˜1.
As shown in this demonstration, doctors may use our framework to learn how much complementary information value the AI models offer over their decisions, and which model offers the most.
@@ 6.2 Deepfake Detection
We study a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022). We select the model in the Deepfake Detection Challenge (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data. We use the human decisions and the human-AI team’s decisions from Groh et al. (2022), who collected judgments on the videos from n=5,524 participants recruited on Prolific. We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from Groh et al. (2022)’s use of mean absolute error, but again we prefer the quadratic score because it is a proper scoring rule where truthfully reporting beliefs maximizes the score. Groh et al. (2022) manually label seven video-level features, which we use as binary indicators in place of Algorithm˜1 in light of the high dimensional signals: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin [^9] The “dark-skin individual” label reflects a subjective visual attribute. It is used here only for robustness analysis—not for causal interpretation or demographic inference.. We show the ACIVs of the seven features over the decisions, the AI predictions, and the human-AI decisions in Figure˜3(b).
@@
We first compare the information value of the AI predictions to that of the human decisions. Figure˜3(b) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$).
@@
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different video-level features over different agents. As shown on the fifth row in Figure˜3(b), the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure˜3(b), the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions. [^10] For space constraints, we refer to the full descriptions and results in Appendices G and F.
@@ 7 Discussion and Limitations
Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al. (2023; 2024)). While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments. Though most of our definitions and analysis are focused on decision tasks with single well-defined payoff function, the framework can be readily extended to more complex decision tasks with an ambiguous set of payoff functions or totally unidentifiable payoff functions. We present a robustness analysis framework in Appendix˜B to calculate the Blackwell ordering of models over proper scoring rules. A demonstration of the robustness analysis is shown in Appendix˜G.
The experiment is designed to provide complementarity by construction in order to function as a proof of concept. We believe future work is needed to evaluate the impact of ILIV-SHAP in real deployment of human-AI decision-making workflows, such as where the human or AI might have additional private information over each other.
8 Reproducibility Statement
The main results of our paper is to provide a framework guiding the analysis of complementary information in human-AI decision-making. We provide a python library in supplementary martial to calculate the quantities defined in our framework.
We provide the data and code (which are all put in separate Jupyter notebooks) to reproduce all the results in our paper, including the demonstrations and the empirical study in the main text, and the observational study and robustness analysis in the Appendix.
@@ Appendix A The Combinatorial Nature of the Value of Signals
When decision-makers are provided with multiple signals, the signals have the combinatorial property by nature. Acknowledgd by recent works in decision theory and game theory (Chen and Waggoner, 2016), one signal may have no information value by itself, but it can be complementary to other signals to provide information value. For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ are uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$. In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff. Though we did not explicitly observe the complementation between signals in our survey of human-AI decision-making tasks (see the results of deepfake detection in supplementary materials), we want to note that our framework can be extended to consider this complementation between signals. We use the Shapley value (Shapley, 1953) to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal. $\phi$ is the average of the marginal contribution of a signal in every combination with other signals.
@@ Appendix B Robust Analysis of Information Order
Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem. However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule. In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). Blackwell’s comparison of signals (Blackwell and others, 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules (Li et al., 2022; Kleinberg et al., 2023).
@@
Proposition˜B.3 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$. Proposition˜B.3 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
@@
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix˜F.
@@ Appendix C Observational Study on the Effect of ACIV on Human-AI Team Performance
In this section, we present the results of an observational study on the effect of $\operatorname{ACIV}$ on human-AI team performance using the dataset from Vodrahalli et al. (2022).
@@
We experiment with the publicly available Human-AI Interactions dataset (Vodrahalli et al., 2022). The dataset comprises 34,783 unique predictions from 1,088 different human participants on four different binary prediction tasks (“Art”, “Sarcasm”, “Cities” and “Census”). In each of the tasks, human participants provide confidence values about their predictions before ($d^{H}$) and after ($d^{H+AI}$) receiving AI advice from a classifier in the form of the classifier’s confidence values ($d^{AI}$). $d^{H}$, $d^{H+AI}$ and $d^{AI}$ are in the range of $[0,1]$.
@@ Results.
We present the $\operatorname{ACIV}$ of the AI confidence values ($d^{AI}$) over human decisions ($d^{H}$) measured by the payoff function as the mean squared error (MSE) in Figure˜4, and show the reduction of the mean squared error (MSE) of the human-AI team over the human-alone baseline in Table˜2. We find that between these four tasks, the $\operatorname{ACIV}$ of the AI over human decisions predicts the improvement of the human-AI team performance over the human-alone baseline.
@@ Appendix D Statistical Tests on Experiment Results
where $s_{1}$ and $s_{2}$ are the standard deviations of the two samples, and $N_{1}$ and $N_{2}$ are the sizes of the two samples (see sample sizes and standard deviations in Table˜5). We used Bonferroni correction as the multiple comparisons correction to control the overall type I error rate. The results are shown in Table˜3 and Table˜4.
Table 3: 95% confidence intervals for the reduction of APE with different AI models and explanations in Figure˜2.
@@
$6.33\%$
$5.90\%,6.76\%]$
$6.22\%$
$5.71\%,6.72\%]$
$5.31\%$
$4.80\%,5.83\%]$
Table 4: P-values for significance tests of reduction in APE between experimental conditions (after Bonferroni correction).
Null Hypothesis ($H_{0}$) P-value
AI1 + ILIV-SHAP + SHAP < AI1 + SHAP 0.002
AI1 + ILIV-SHAP + SHAP < AI1 + No explanation 0.010
AI1 < AI2 0.555
AI1 + ILIV-SHAP + SHAP < AI2 + ILIV-SHAP + SHAP <0.001
@@
Appendix E Cognitive Loads and Anchoring Effects in Experiment
We also examine the spent time on task and the degree of anchoring on the human versus AI decisions by looking at the distance between human-AI team decisions and AI or human-alone decisions.
Time spent.
Table˜6 shows that there is no significant increase in the duration of the experiment for ILIV-SHAP (23.3 [13.1, 32.1] minutes for AI1 with ILIV-SHAP and 22.2 [11.9, 28.0] minutes for AI2 with ILIV-SHAP, where the square brackets show the 25% and 75% quantiles).
Anchoring effects.
Table˜6 also shows that ILIV-SHAP does not increase anchoring on the AI or human alone decisions, while the presence of SHAP alone tends to make the participants anchor more on the AI model. With no explanation, the participants anchor more on their own decisions in the first round.
Table 6: The number of participants (after filtering based on the criteria in the pre-registration), the mean duration of the experiment, and the anchoring on AI and human for each condition. Anchoring on AI is calculated as the mean of the absolute difference between the human-AI team’s prediction and the AI’s prediction normalized by the actual sale price, $|d^{\text{Human-AI}}-d^{\text{AI}}|/\omega$, and anchoring on human is calculated as the mean of the absolute difference between the human-AI team’s prediction and the human’s prediction normalized by the actual sale price, $|d^{\text{Human-AI}}-d^{\text{Human}}|/\omega$. The square brackets indicate the 25th and 75th percentiles.
Condition # of participants Mean Duration (minutes) Anchoring on AI ($\downarrow$) Anchoring on Human ($\downarrow$)
AI1 + ILIV-SHAP+SHAP 70 23.3 [13.1, 32.1] 0.303 [0.0277, 0.253] 0.404 [0.0746, 0.491]
AI1 + SHAP 79 21.3 [12.0, 27.9] 0.211 [0.0201, 0.222] 0.427 [0.0721, 0.533]
AI1 + No Explanation 69 22.4 [14.5, 28.4] 0.226 [0.0276, 0.258] 0.367 [0.0740, 0.483]
AI2 + ILIV-SHAP+SHAP 67 22.2 [11.9, 28.0] 0.206 [0.0273, 0.260] 0.458 [0.0840, 0.613]
AI2 + SHAP 62 23.1 [13.2, 27.0] 0.203 [0.0215, 0.224] 0.456 [0.0777, 0.602]
AI2 + No Explanation 74 22.5 [12.7, 30.5] 0.259 [0.0265, 0.260] 0.434 [0.0611, 0.501]
Appendix F Demonstration I: Model Comparison on Chest Radiograph Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025). We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
@@ Appendix F Demonstration I: Model Comparison on Chest Radiograph Diagnosis
F.1 Data and Model
We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023). The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. (2019) and Heidenreich et al. (2022). We label the radiology reports by a rule-based tool (Irvin et al., 2019) and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports. The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$. We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017). Our training set contains 12,228 images, and the validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models is: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, and VisionTransformer with $0.80$.